OpenStudy (anonymous):

open parentheses 5 plus square root of 11 close parentheses open parentheses 5 minus square root of 11

OpenStudy (shamim):

\[(5+\sqrt{11})(5-\sqrt{11})=14\]

OpenStudy (shamim):

is it wt u wanted

OpenStudy (anonymous):

yes thanks sharmin

OpenStudy (shamim):

u r most welcome

OpenStudy (shamim):

u can give me a medal

OpenStudy (anonymous):

sharmin i got a new problem could you help me out& a medal im really new to this site so idk what that means?

OpenStudy (shamim):

ok u can tell me ur new problem

OpenStudy (anonymous):

square root of 12 open parentheses square root of 3 end root plus 2 close parentheses plus square root of 2 open parentheses square root of 8 end root minus square root of 6

OpenStudy (shamim):

i m shamim @daze

OpenStudy (shamim):

\[\sqrt{12}(\sqrt{3}+2)+\sqrt{2}(\sqrt{8}-\sqrt{6})\]

OpenStudy (shamim):

\[or, \sqrt{36}+2\sqrt{12}+\sqrt{16}-\sqrt{12}\]

OpenStudy (shamim):

\[=6+\sqrt{12}+4\]

OpenStudy (shamim):

\[=10+\sqrt{12}\]

OpenStudy (shamim):

\[10+2\sqrt{3}\]

OpenStudy (shamim):

is it u expected frm me